How many grams of glucose, Co H12 O6 would you need to prepare 2.0 1 of 2 % glucose solution?


1
Expert's answer
2021-04-27T07:50:11-0400

2% of (2/100× 100)g = 2g glucose solutions in 100ml is solution

Thus 200ml = 200×(2/100) = 4g of glucose